OpenStudy (anonymous):

what is the equation of the line that has a slope -2 and passes through the point (3,6)

OpenStudy (anonymous):

the equation of a line is normally y=mx+b where m= the slope so y=-2x+b and b is the y intercept. you can find b by plugging in the point given (3,6) into what you already have so 6=-2(3)+b and solve. after you have b put it all together :0

OpenStudy (anonymous):

Its been awhile since I've done algebra like that but Im 99% sure that's what you do it

OpenStudy (anonymous):

makes sense to me xlovely_Lizardx.......now, 2479hann....: 6 = -2(3) + b -- solve for b can you do that ?